Ace Industrial Machines issued 190,000 zero coupon bonds 7 years ago. The bonds originally had 30 years to maturity with a yield to maturity of 6.6 percent. Interest rates have recently decreased, and the bonds now have a yield to maturity of 5.7 percent. The bonds have a par value of $2,000. If the company has a $88.2 million market value of equity, what weight should it use for debt when calculating the cost of capital? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., .1616.)
Answer #1
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =23
Bond Price =∑ [(0*2000/100)/(1 + 5.7/100)^k]     +   2000/(1 + 5.7/100)^23
                   k=1
Bond Price = 558.86

MV of bond = price*bonds = 558.86*190000

=

106183400
Total Capital value = Value of Debt + Value of Equity
=106183400+88200000
=194383400
Weight of Debt = Value of Debt/Total Capital Value
= 106183400/194383400
=0.5463